OpenStudy (anonymous):

pic a number dont tell me multypluy it by 2 add 6 divide by 2 subtract the number from the start

OpenStudy (bahrom7893):

[(2x+6)/2]-x = x + 3 - x = 3, so the answer will be 3 no matter what.

OpenStudy (anonymous):

Aw, you ruined the magic...

OpenStudy (bahrom7893):

haha =) Math is magic!